Report of Activities on the RVIB N.B. Palmer Cruise 02-02

12 and 13 April 2002

 

The transit south from Punta Arenas, Chile to our survey grid on western Antarctic Peninsula continued for a fourth and fifth day. The early morning hours of the 12th of April found the Palmer in rough seas and winds hovering about 30 kts and still out of the west southwest (250). We passed the Argentine 200 mile limit in the wee hours and began to make some scientific observations. There was a noticeable drop in both the sea (1.7C) and air temperature (1.2C) about the time we left the Argentine economic zone which marked our crossing of the polar front. By mid-day the winds were dropping and the seas moderating. Late in the afternoon, the winds died down to the 17 to 21 kt range out of the west-northwest (300). The barometer was still above a 1000 (1001.0) and the air temperature (1.6C) was colder than the seawater (2.06C). The clouds remained along with a light drizzle. Low visibility made it hard for the bird and marine mammal observers to conduct their surveys.

 

A science meeting was held at 1300 on the 12th and the different scientific parties on board reviewed their scientific objectives and outlined what they planned to do at the various stations. There was consensus that a test station some distance from the first station in the grid was needed and that was programmed into the schedule.

 

Just after sunrise (~0800) on the morning of the 13th , the test station began about 100 nm north of Grid Station 1. The sea surface was almost glassy and only a low swell was running. Fog hung over the sea surface, but it was not so thick that the ship needed to slow from its 11 knot pace in reaching the test station. But the skies were a hazy light blue above and the winds light. The air temperature (-1.9C) and sea temperature (-0.03C) continued to decline. The CTD was quickly deployed. The profile to 500 meters went well and except for a couple of bottles that did not close properly, the cast was successful. This was followed by a BIOMAPER-II deployment to 180 m, an Acoustic Properties of Plankton measurement system deployment, and a MOCNESS tow. A second deployment of BIOMAPER-II late in the afternoon was needed to fine tune the towing configuration.

 

With the completion of the test station, we again set sail for Station 1. Sea conditions changed significantly during the day. By noon it was overcast, but the sun still shone through a bit. Late in the afternoon, it was sleeting lightly and the wind had picked up. Now at 2153 on the 13th, the winds are back up around the 30 kt mark out of the west (274). The barometer has been falling some and is now at 983.7 mlb. Air temperature is just above freezing (0.8C) and the water temperature just below (-0.036C). Our current position is -64 32.125S; -69 45.393W and we are steaming hard (11 kts) to our first station #1.

 

CTD Group report (John Klinck, Tim Boyer, Chris Mackay, Julian Ashford, Andres Sepulveda, Kristin Cobb)

The CTD group started sampling early in the morning on April 11 as the Palmer sailed out of the Argentine EEZ with an XBT cast to 760 meters. For the next two days (ending in the morning of 13 April), 38 XBT casts were made to 760 m depth at a station spacing of 10 nm. Of these 38 casts, one probe failed on load, and four more failed due to spikes or erratic temperatures. The track followed 65W longitude for about a third of the way across the passage, then turned southwest towards station 1.

 

A section plot of temperature from these data reveal that the polar front was located between the first two stations, just south of the 200 mile limit.

 

South of this for the next 200 km, the surface waters ranged from 1 to 2C with a thickness of about 80 m. Below the surface layer was a clear Winter Water (WW) layer about 100 m thick with minimum temperatures from -0.5 to -1.0C. Curiously, one station showed a minimum temperature above +0.5 indicating an almost complete erosion of the WW.

 

In the southern 200 km of the section, surface waters were much colder (-0.5 to 0.5C) with a thickness of about 80 m. The WW was thinner (about 50 m) with temperatures below -1.0C. The depth of base of the WW layer shifted by as much as 100 m over distances of 50 km or so. As we moved southward, the surface layer becomes cooler and merged with the WW. From the section as a whole, the base of the WW layer rises from 250 m in the north to about 100 m in the south.

 

At the southernmost cast, the mixed layer is rather warm (0.0C), 100 m thick with no trace of the WW. The southern 3 casts show almost no subsurface temperature minimum (WW).

 

We did a test cast of the CTD at 0830 on 13 April. The low winds and calm sea made it a pleasure. The cast was to 500 m with all sensors on, including the FRRF, but not the microstucture probe. The brackets are still being worked on and we want a minimum of attaching and detaching of this sensor. Due to the deep cast at station 1, we decided to wait until station 2 to use this new sensor. With success, we could leave it on the rosette until the next deep station (12).

 

The cast was accomplished with the expected delays and confusion. In the end, the cast was completed successfully. A uniform mixed layer was observed to about 75 m with a thin (15 m) only slightly cooler (-0.5C) WW layer. The flourometer showed a strong response (0.2 mg/l) in the surface layer and instrument noise below. The new oxygen sensor adjusted rapidly and displayed no erratic behavior.

 

As a test of bottle integrity, all (23) bottles were closed at 500 m. Three bottles did not trip. We drew 20 oxygen samples as practice and to start working with the oxygen titrator. We also drew 8 salinity samples to start working with the salinometer.

 

Several problems with software were detected which are being worked on. These have to do with processing and exporting the data as well as printing the cast plot. It has been a busy two days, with busier days to come.

 

Marine Mammal report (Debra Glasgow)

This report summaries observations since leaving Punta Arenas Chile. After departing Punta Arenas, Chile at 1100 on 9 April, the Cajun Cruncher was picked up around 1145. While there, 3 unidentified dolphins were observed by Eric Chapman to come within 10 meters of the bow. Some white and grey markings were seen on the animals flanks. They came directly towards the ship, but were seen only briefly before disappearing. At 1610 in the Straits of Magellan, a possible Commersons dolphin, which surfaced about 7 meters from the bow, was seen by Matt Becker. During the period that we were towing the BIOMAPER-II, an incidental watch was kept from 1600 until dark (1900), but there were no further sightings.

 

On 10 April, at sunrise 0700), conditions for cetacean survey were not good with the wind speed at 35-40 knots from 220-230 continuing throughout the day. The increasing swell throughout the day and the worsening weather conditions precluded any decent cetacean survey.

 

No watch was kept on 11 April due to high wind and seas until the weather calmed somewhat to 30 knots around 1600 and the ship picked up speed to 9.4 knots. In spite of a haze on the horizon and occasional sleet squalls, an incidental watch was kept on the bridge from 1600 until stopped by encroaching darkness and poor visibility at 1745. No cetaceans were sighted.

 

An overcast and generally foggy day today especially in the afternoon. Three and a half hours full survey was achieved on the morning of 12 April, though no cetaceans were sighted by the observer. Around 10 a.m. Yulie Serebrennikova sighted 1 black and white dolphin (estimated about 3 meters in size) porpoising briefly by the port bow. About the same time Ana Sirovic launched a sonobuoy. After the science meeting in the afternoon, a full survey was attempted and then downgraded to an incidental watch from the bridge for 2 hours 20 minutes until dark as the persistent fog limited visibility to mostly within 300 meters. At 1545 and 1627 porpoising penguins were sighted.

 

An all day incidental watch was conducted on 13 April beginning at 0730 mainly from the ice tower through variable visibility conditions. The windows in the ice tower this morning were covered in ice inside and out and some of the morning's observations were from the bridge while the windows thawed. Though the wind and sea were slight for most of the day, fog kept visibility low for most of the time with rare clear patches. During one clear time, one unidentified large whale was sighted at 1247, but proved difficult to track and identify as we were often changing direction while we tested various instruments. The ship traveled at 1-3 knots throughout the day and because of this, the fog, and the constant changes in direction, a full survey mode was not possible. Incidental mode observing ceased at 1600 due to rain and fog limiting visibility to less than 200 meters. Large numbers of birds accompanied the ship throughout the day, increasingly so as the day wore on.

 

Marine Mammal passive listening report (Ana Sirovic)

After steaming out of the Argentine 200 mile EEZ, a total of 3 sonobuoys were deployed on 12 and 13 April. Blue whale calls were heard on one sonobuoy and possible fin whale calls were heard on another. The blue whale appeared to have been swimming north of the sonobuoy deployment location. The system seems to be working well and antenna range is satisfactory.

 

Sea Birds (Erik Chapman)

On 13 April, incidental observations were made from the bridge while the BIOMAPER-II and MOCNESS were being towed at the test station. There were about 170 to 200 birds around the ship, about 70% of which were Cape Petrels and the rest were mainly Southern Fulmars. Cape Petrels were actively surface-seizing prey in the prop wash behind the ship and anywhere turbulence caused by the ship appeared to be bringing water up to the surface. The birds were probably feeding on larval krill and copepods that were found in the surface nets during the MOCNESS tow. We also saw a few Grey Headed Albatross, Southern Giant Petrels and Blue Petrels, and an unidentified Storm Petrel in the same area. Hopefully, the real data collection will begin tomorrow!

 

Zooplankton (MOCNESS/BIOMAPER-II) report (Carin Ashjian, Peter Wiebe)

The BIOMAPER-II test involved a towyo in which the towed body was sent to 180 meters and then hauled back towards the surface while making measurements with the five frequency acoustic system, the two camera video plankton recorder, and the CTD, fluorometer, and transmissometer sensors. On the return to the surface, the fish was held at 80 meters and a noise test of the acoustic system was conducted. All three sensor system performed well. But the towed body performance was poor and adjustments to optimize its roll, pitch and yaw characteristics were needed. During the first part of the tow, significant numbers of krill larvae were observed near the surface in the VPR images and then later in the tow, several large patches of scatterers, probably krill, were centered about 180 meters below the surface. In the pycnocline below the Winter Water layer at the test station, the echograms of the higher frequencies (200 and 420 kHz) displayed horizontal layers of higher backscattering separated by lower backscattering that appeared to be associated with the physical structure of that portion of the water column. The microstructure sensor when used on the CTD should help determine this. A second deployment of BIOMAPER-II took place around 1700 after making some adjustments to tail elevator to see if the towing performance was improved.

 

The first MOCNESS tow of the cruise was conducted at the test station on 13 April. The short tow to 200 m, sampling at 25 m intervals, went very well except for the lack of net response signals; we will adjust the net response for the next tow. Many salps were collected in the 100-200 m depth intervals. Amphipods were collected in the 200-175 m depth range. Chaetognaths were present from 125-75 m. Krill furcilia were present from 100-25 m, along with copepods. Only copepods were observed in the upper 25 m; they appeared to be genus Calanus and were much more abundant at this depth than elsewhere (even though we also filtered more water with this net). The water column oblique net was preserved in ethanol for genetic analysis.

 

Cheers, Peter